OpenStudy (anonymous):

compute limit [x->0] (sqrt(cos(x)) - (cos(x))^(1/3))/(sin(x)^2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ \sqrt{\cos(x)} - \sqrt[3]{\cos(x)} }{ \sin^2(x) }\]

OpenStudy (kainui):

Are you allowed to use L'Hopital's rule?

OpenStudy (anonymous):

no :(

OpenStudy (kainui):

That sucks. I'm trying to figure it out without it then haha. Give me a second, I am trying to multiply it by [cos^(1/2)x+cos^(1/3)x]/[cos^(1/2)x+cos^(1/3)x] right now.

OpenStudy (kainui):

Actually before that, maybe you should try splitting the limit up into two parts and solving those individually.

OpenStudy (kainui):

What sort of tricks are you allowed to use if not L'H rule?

OpenStudy (anonymous):

Every other tricks are allowed ;)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sqrt{\cos x}-\left( \cos x \right)^{\frac{ 1 }{ 3 }} }{\sin ^{2}x }\] \[=\lim_{x \rightarrow 0}\frac{ \sqrt{\cos x}-1+1-\left( \cos x \right)^{\frac{ 1 }{ 3 }} }{ \sin ^{2}x }\] \[=\lim_{x \rightarrow 0}\frac{ \sqrt{\cos x}-1 }{\sin ^{2}x }+\lim_{x \rightarrow 0}\frac{ 1-\left( \cos x \right)^{\frac{ 1 }{ 3 }} }{ \sin ^{2 }x}=L1+L2\] for L1 multiply the numerator and denominator by \[\sqrt{\cos x}+1 and solve\]

OpenStudy (anonymous):

\[1+\left( \cos x \right)^{\frac{ 1 }{ 3 }}+\left( \cos x \right)^{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

multiply the numerator and denominator of L2 as written above.